# Persistent Homology for High-dimensional Data Based on Spectral Methods

### v2

@@ Appendix B Noise in high-dimensional spaces eventually dominates structure
Appendix C Effective resistance as a spectral method
@@
Let $G$ be a weighted, connected graph of $n$ nodes. Denote by $A=(a_{ij})_{i,j=1,\dots,n}$ the weighted adjacency matrix whose entries $a_{ij}=a_{ji}$ equal the edge weight of edge $ij$ if edge $ij$ is part of the graph and zero otherwise. Further denote by $D=\text{diag}(d_{i})$ the degree matrix, where $d_{i}=\sum_{j=1,\dots,n}a_{ij}$ are the degrees. We define $\text{vol}(G)=\sum_{i}d_{i}$. Let further $A^{\text{sym}}=D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$ and $L^{\text{sym}}=I-A^{\text{sym}}$ be the symmetrically normalized adjacency matrix and the symmetrically normalized graph Laplacian. Denote the eigenvectors of $L^{\text{sym}}$ by $u_{1},\dots,u_{n}$ and their eigenvalues by $\mu_{1},\dots,\mu_{n}$ in increasing order. The eigenvectors and eigenvalues of $A^{\text{sym}}$ are $u_{1},\dots,u_{n}$ and $1-\mu_{1},\dots,1-\mu_{n}$.
Definition C.1.
@@
$\frac{1}{\text{vol}(G)}H_{ij}=\frac{1}{d_{j}}+\langle b_{j},A^{\text{sym}}(b_{j}-b_{i})\rangle+\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle A^{\text{sym}}b_{j},u_{r}(u_{r})^{\top}(A^{\text{sym}}(b_{j}-b_{i}))\rangle,$ (20)
where $b_{i}=\frac{1}{\sqrt{d_{i}}}\hat{e}_{i}=D^{-\frac{1}{2}}\hat{e}_{i}$ with $\hat{e}_{i}$ the $i$-th standard basis vector.
@@
$\quad+\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle A^{\text{sym}}b_{j},u_{r}(u_{r})^{\top}(A^{\text{sym}}(b_{j}-b_{i}))\rangle$ (23)
$\quad+\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle A^{\text{sym}}b_{i},u_{r}(u_{r})^{\top}(A^{\text{sym}}(b_{i}-b_{j}))\rangle$ (24)
@@
$\quad+\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle A^{\text{sym}}(b_{j}-b_{i}),u_{r}(u_{r})^{\top}(A^{\text{sym}}(b_{j}-b_{i}))\rangle$ (26)
For ease of exposition, we treat both summands separately. By unpacking the definitions and symmerty of $D$, we get
@@
$\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle A^{\text{sym}}(b_{j}-b_{i}),u_{r}(u_{r})^{\top}(A^{\text{sym}}(b_{j}-b_{i}))\rangle$ (30)
$=$ $\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle(b_{j}-b_{i}),(A^{\text{sym}}u_{r})(A^{\text{sym}}u_{r})^{\top}(b_{j}-b_{i}))\rangle$ (31)
$=$ $\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle(b_{j}-b_{i}),(1-\mu_{r})u_{r}((1-\mu_{r})u_{r})^{\top}(b_{j}-b_{i}))\rangle$ (32)
@@
The corrected version of effective resistance and diffusion distances are, in general, not proper metrics, unlike the naive effective resistance [27, Corollary 2.4.].
Proposition C.3.
Neither corrected effective resistance nor diffusion distances are necessarily positive on distinct points. Moreover, corrected effective resistance does not always satisfy the triangle inequality.
@@
Consider the unweighted chain graph with three nodes (Figure S2 left). The uncorrected effective resistance between the first and last node is $2$. So the corrected effective resistance between these distinct nodes is $2-1-1=0$. On the same graph, the random walker is necessarily at node $1$ after one step, independent whether it started at node 00 or $2$, so that the diffusion distance with $t=1$ between two distinct nodes is zero.
Consider now an unweighted graph with $5$ nodes, the first three of which form a triangle and the other two a chain connected to the triangle (Figure S2 right). Then
@@
So that the triangle $0,2,4$ violates the triangle inequality
@@
Nevertheless, using the square root of corrected effective resistance (which does satisfy the triangle inequality) as input to persistent homology amounts to taking the square root of all birth and death times found with the corrected effective resistance as computing persistent homology commutes with strictly monotonic maps. This maintains the excellent loop detection performance of corrected effective resistance (Figure S16). We therefore believe that it is not a problem for our persistent homology application that corrected effective resistance may fail to satisfy the triangle inequality.
@@
Figure S2: Counterexample graphs for Proposition C.3
Appendix D Effective resistance integrates diffusion distances
First, we describe a connection between diffusion distances and naive effective resistance communicated to us by Mémoli [46]. To the best of our abilities, we could not find another reference. We use the same notation as in Section 6 and Appendix C.
Below, we will speak of diffusion distances at half-integer time points, although the random walk analogy does not extend to half-steps. What we mean is inserting a half-integer value for $t$ in Eq. (4).
Proposition D.1.

### v3

@@ Appendix B Noise in high-dimensional spaces eventually dominates structure
Appendix C Persistent homology with PCA
Hiraoka et al. [42] recommend combating the curse of dimensionality by performing a normalized PCA before applying persistent homology. We explore this approach here in our setting.
Let $X=(x_{1},\dots,x_{n})^{T}\in\mathbb{R}^{n\times d}$ be the centered data matrix with each data point as a row. Let $X=USV^{T}$ be the singular value decomposition of $X$ with $U$ and $V$ orthogonal matrices and $S\in\mathbb{R}^{n\times d}$ diagonal with decreasing non-negative values on the diagonal. PCA reduces the dimensionality to $k\leq d$ dimensions by considering the first $k$ columns of $US$, while normalized PCA instead considers the first $k$ columns of $U$.
We ran experiments on the 1D datasets for both versions with varying number of principal components (PCs) (Figure S2). Normalization provided no consistent benefit. The performance was worse for higher number of PCs as more noise remained after PCA. Using fewer PCs than the dataset’s true dimensionality (Figure S2f, j) also led to poor performance. Knowing the true dimensionality of the data is therefore crucial for this approach, but true dimensionality is typically not available in real-world applications. Effective resistance outperformed PCA preprocessing on the linked circles and eyeglasses dataset and showed similar performance on the circle.
Figure S2: Loop detection score for effective resistance and PCA on three 1D toy datasets.
Appendix D Continuity of the hole detection score
In this section we prove that our hole detection score is a continuous map, as long as the persistence diagrams have sufficiently many points. This means that small changes in the persistence diagram result in small changes of the hole detection score, a desirable property for a performance measure.
To state our result formally, we recall some definitions from Cohen-Steiner et al. [19]:
Definition D.1 (Multiset).
A multiset is a set in which each element has a multiplicity in $\mathbb{N}_{>0}\cup{\infty}$.
Definition D.2 (Persistence diagram).
A persistence diagram is a multiset with elements from $\{(b,d)\in(\mathbb{R}\cup\{-\infty,\infty\})^{2}\mid b<d\}$ counted with multiplicities, together with the diagonal $\{(b,b)\in(\mathbb{R}\cup\{-\infty,\infty\})^{2}\}$ counted with infinite multiplicity.
Since we build persistence diagrams only for distances-based Vietoris–Rips complexes and do not consider 00-dimensional homologies in this paper, no birth times will be $\infty$- ∞. Moreover, since we do not consider infinite distances (Appendix I), no death times will be infinite either. Therefore, in the following, we will treat persistence diagrams as multisets with elements in $\mathbb{R}^{2}$.
Definition D.3 (Bottleneck distance).
Let $D$ and $D^{\prime}$ be persistence diagrams. Then the bottleneck distance between them is given by
$d_{B}(D,D^{\prime})=\inf_{\eta:D\to D^{\prime}}\sup_{x\in D}\|x-\eta(x)\|_{\infty},$
where the infimum is over the bijections between the multisets $D$ and $D^{\prime}$.
A bijection $\eta$ realizing the bottleneck distance between persistence diagrams has three types of paired points: Pairs with both points off the diagonal, pairs with exactly one point off the diagonal, and pairs with both points equal and on the diagonal. Intuitively, the bottleneck distance matches the off-diagonal points between each other. However, the second type of pairs is needed in case the persistence diagrams do not have the same number of points off the diagonal. This is why the full diagonal with infinite multiplicity is considered to belong to each persistence diagram. The third type of pairs does not contribute to the bottleneck distance.
We can now proceed with our proofs. Let $\mathcal{D}$ be the space of all persistence diagrams with finitely many points off the diagonal, endowed with the topology induced by the bottleneck distance $d_{B}$.
Lemma D.4.
For any $m\in\mathbb{N}$, the map $p_{m}:\mathcal{D}\to\mathbb{R}_{\geq 0}$ assigning to a diagram its $m$-th largest persistence value is continuous.
Proof.
For a persistence diagram $D\in\mathcal{D}$ and a feature $f=(\tau_{b},\tau_{d})\in D$ we denote by $p(f)=\tau_{d}-\tau_{b}$ the persistence of $f$. Note that for two features $f,f^{\prime}$ in a persistence diagram, we have $|p(f)-p(f^{\prime})|\leq\sqrt{2}\|f-f^{\prime}\|_{\infty}$.
We need to show that for any $\varepsilon>0$ and any $D\in\mathcal{D}$, we can choose some $\delta>0$ such that $|p_{m}(D^{\prime})-p_{m}(D)|<\varepsilon$ for all $D^{\prime}$ whose bottleneck distance from $D$ is at most $\delta$, i.e., $D^{\prime}\in B_{\delta}(D)$. Recall that in this setting, by definition of the bottleneck distance, there is a bijection (of multisets) $\eta:D\to D^{\prime}$ such that for all $f\in D$ we have $\|f-\eta(f)\|_{\infty}<\delta$.
Continuity of $p_{m}$ is easier to demonstrate at diagrams $D\in\mathcal{D}$ for which all off-diagonal points have different persistences. We discuss this case first. Assume $D$ has $l$ off-diagonal points. Denote by $f_{k}$ the feature with $k$-th largest persistence.
Let $\tilde{\delta}=\min\{p(f_{k})-p(f_{k+1})\mid k\leq l\}$ and $\delta=\min\big{(}\tilde{\delta}/(2\sqrt{2}),\varepsilon/(2\sqrt{2})\big{)}$. Consider $D^{\prime}\in B_{\delta}(D)$. By the choice of $\delta$, we have that $|p(\eta(f_{k}))-p(f_{k})|<\tilde{\delta}/2$, so that the values $p(\eta(f_{1})),\dots,p(\eta(f_{l})),0$ are all distinct and strictly decreasing. As a result, for all $m\leq l$ we have $p_{m}(D^{\prime})=p(\eta(f_{m}))$ and thus
$|p_{m}(D^{\prime})-p_{m}(D)|=|p(\eta(f_{m}))-p(f_{m})|<\sqrt{2}\delta<\varepsilon.$
For $m>l$, we have $p_{m}(D)=0$ and any $f^{\prime}\in D^{\prime}$ with $p_{m}(D^{\prime})=p(f^{\prime})$ is paired with a point on the diagonal of $D$ under $\eta$. Hence
$|p_{m}(D^{\prime})-p_{m}(D)|=|p(f^{\prime})-0|<\sqrt{2}\delta<\varepsilon.$
The case where $D$ might contain off-diagonal points with identical persistence follows the same overall argument, but requires more careful bookkeeping. Assume $D$ has $l$ off-diagonal points. Let $\tilde{\delta}=\min\big{(}\{p_{k}(D)-p_{k+1}(D)\mid k\leq l\}\backslash\{0\}\big{)}$. Let
$\delta=\min\big{(}\tilde{\delta}/(2\sqrt{2}),\varepsilon/(2\sqrt{2})\big{)}.$
Set
$F_{k}=\{f\in D\mid p(f)=p_{k}(D)\}$
for $k\leq l$ and $F_{l+1}$ equal to the diagonal. These multisets all contain at least one element. We have $p_{k}(D)=p_{k+1}(D)$ if and only if $F_{k}=F_{k+1}$.
As a result, for all $k\leq l$, we have $|\bigcup_{i=1}^{k}F_{i}|=k$ if and only if $F_{k}\neq F_{k+1}$ which in turn holds if and only if $p_{k}(D)\neq p_{k+1}(D)$.
For any $f,\tilde{f}\in D$ with $p(f)>p(\tilde{f})$ we have
$p(\eta(f))-p(\eta(\tilde{f}))$ $=p(\eta(f))-p(f)+p(f)-p(\tilde{f})+p(\tilde{f})-p(\eta(\tilde{f}))$
$\geq p(f)-p(\tilde{f})-|p(\eta(f))-p(f)|-|p(\tilde{f})-p(\eta(\tilde{f}))|$
$>p(f)-p(\tilde{f})-\tilde{\delta}$
$\geq 0,$ (18)
so $p(\eta(f))>p(\eta(\tilde{f}))$. We will now show that if $m\leq l$ and $f^{\prime}\in D^{\prime}$ such that $p_{m}(D^{\prime})=p(f^{\prime})$, then $\eta^{-1}(f^{\prime})\in F_{m}$.
Let $a$ be the highest number below $m$ such that $F_{a}\neq F_{m}$. Then
$\left|\bigcup_{i=1}^{a}\eta(F_{i})\right|=\left|\bigcup_{i=1}^{a}F_{i}\right|=a.$
and $\bigcup_{i=1}^{a}\eta(F_{i})$ contains the $a<m$ features of $D^{\prime}$ with smallest persistence. In particular, $\eta^{-1}(f^{\prime})\notin\bigcup_{i=1}^{a}F_{i}$.
Let $b$ be the largest number such that $F_{h}=F_{m}$. By a similar argument, $\bigcup_{i=1}^{b}\eta(F_{i})$ contains the $b\geq m$ features in $D^{\prime}$ with largest persistence, including $f^{\prime}$. Thus, $\eta^{-1}(f^{\prime})\in\bigcup_{i=a+1}^{b}F_{i}$. But by choice of $a$ and $b$, we have $F_{a+1}=\dots=F_{m}=\dots=F_{b}$, so $\eta^{-1}(f^{\prime})\in F_{m}$.
Thus,
$|p_{m}(D^{\prime})-p_{m}(D)|=|p(f^{\prime})-p(\eta^{-1}(f^{\prime}))|\leq\sqrt{2}\|f^{\prime}-\eta^{-1}(f^{\prime})\|_{\infty}<\varepsilon.$
For $m>l$, we have $p_{m}(D)=0$ and $F_{m}$ is the set of points on the diagonal of $D$. Moreover,
$\left|\bigcup_{i=1}^{l}F_{i}\right|=l=\left|\bigcup_{i=1}^{l}\eta(F_{i})\right|,$
so that any $f^{\prime}\in D^{\prime}$ with $p_{m}(D^{\prime})=p(f^{\prime})$ must be in the image of the diagonal under $\eta$ and we can proceed as in the case of unique positive persistences. ∎
Our hole detection score is a continuous score if there are at least $m$ points in the persistence diagram, and not continuous otherwise:
Proposition D.5.
Let $\mathcal{D}$ be the space of all persistence diagrams with finitely many points off the diagonal, endowed with the topology induced by the bottleneck distance $d_{B}$. Let $m\in\mathbb{N}$ and define $\mathcal{D}^{\prime}$ as the subset of $\mathcal{D}$ with at least $m$ points off the diagonal. Then the map $s_{m}:\mathcal{D}\to\mathbb{R}_{\geq 0}$ is continuous at all diagrams in $\mathcal{D^{\prime}}$ but is discontinuous at all diagrams in $\mathcal{D}\backslash\mathcal{D^{\prime}}$.
Proof.
The discontinuity of $s_{m}$ at diagrams with less than $m$ points happens when adding the $m$-th point to a diagram. Let $D$ be a persistence diagram with $l<m$ points off the diagonal. Let $\varepsilon>0$ and define $D^{\prime}$ as the persistence diagram obtained by adding $m-l$ copies of $(0,\varepsilon/2)$. Then $s_{m}(D)=0$ and $s_{m}(D^{\prime})=1$, because $D^{\prime}$ has exactly $m$ off-diagonal points, but $d_{B}(D,D^{\prime})<\varepsilon$. Thus, $s_{m}$ is not continuous at $D$.
For the continuity part, we write $s_{m}$ as the composition of two continuous maps. The map $q:\mathbb{R}_{>0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0},$ $(x,y)\mapsto(x-y)/x$ is continuous. Furthermore, since each diagram in $\mathcal{D}^{\prime}$ has at least $m$ points off the diagonal, $p_{m}(D)\in\mathbb{R}_{>0}$ for all $D\in\mathcal{D}^{\prime}$. By lemma D.4 the map $p_{m}\times p_{m+1}:\mathcal{D}^{\prime}\to\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}$ is continuous. Finally, $s_{m}:\mathcal{D}^{\prime}\to\mathbb{R}_{\geq 0}$ factors as $q\circ(p_{m}\times p_{m+1})$, showing its continuity. ∎
In other words, on persistence diagrams with sufficiently many points our hole detection score is continuous. This makes our score robust against small perturbations in the persistence diagram, a property expected for a reliable score.
In the typical case, where the persistence diagram contains a noise cloud of many points close to the diagonal and, possibly, some outliers far away from the diagonal, the requirement on the number of points is satisfied. Only in the case where we expect a large gap after the $m$-th feature, but the diagram does not even have $m$ features, can there be a sudden jump in our hole detection score.
Appendix E Alternative performance scores
While our hole detection score corresponds to an intuitive visual assessment of persistence diagrams and is continuous (Appendix D), in this section we discuss two alternative performance scores. We argue that both of them are less suited for our benchmark.
E.1 Widest gap score
A natural way to interpret a persistence diagram is to deem all features above the widest gap in persistence values as true features and the rest as noise, i.e., infer that a diagram has $a$ true features if $a=\text{argmax}_{\alpha}(p_{\alpha}-p_{\alpha+1})$.
We considered the binary score which is equal to 1 if the number of features above the widest gap equals the number of ground-truth features and 0 otherwise. We call it the widest gap score.
Note that our hole detection score can be lower than 1 even if the persistence diagram clearly shows the correct number of outlier features (e.g. $s_{m}\approx 0.5$ for $\sigma=0.2$ in Figure 3b). In contrast, the widest gap score is equal to 1 in this case.
On the other hand, the binary nature of the widest gap score leads to instability in case of nearly equisized gaps. For instance in Figure 9d, a small perturbation could make the gap after the most persistent feature the widest, dropping the widest gap score from 1 to 0. In contrast, our $s_{2}\approx 0.9$ correctly reflects the two clear outlier features.
We evaluated several distances on the three 1D datasets using the widest gap score in Figure S3. The binary nature of the widest gap score led to high-variance results even though we increased the number of random seeds from 3 to 10. Spectral methods still performed better than Euclidean, Fermat, and DTM distances. However, the widest gap score led to some false positives. For example, for the Fermat distance and $\sigma=0.3$ for the circle dataset, in 9 of 10 trials the widest gap appeared after the first feature. However, all 9 persistence diagrams looked like noise clouds and in 8 cases the most persistent feature’s representative cycle was clearly a noise feature (Figure S3d – e). Conversely, for the effective resistance there was always a clear outlier in the diagram and the representative did follow the ground-truth loop (Figure S3f – g). Although the Fermat distance failed in this setting and the effective resistance performed well, their widest gap scores were both high.
For these reasons, we prefer our hole detection score to the widest gap score.
Figure S3: Top row: Widest gap score on the three 1D toy datasets. We used 10 random seeds here. Spectral methods outperformed Fermat, DTM, and Euclidean in this score. Bottom row: Example persistence diagrams and representatives for the toy circle with $\sigma=0.3$ with Fermat distance and effective resistance illustrate that despite the high score for both, Fermat distance actually failed.
E.2 Cycle matching
The need to distinguish between true signal and noise features in a persistence diagram motivated Reani and Bobrowski [64] to develop cycle matching, a technique that quantifies the correspondence of features in two persistence diagrams by a prevalence score in $[0,1]$. True signal corresponds to features that reappear with high prevalence in variations of the data obtained, e.g., via resampling.
This approach to detecting the true topology of high-dimensional data is orthogonal to our exploration of different input distances. Indeed, cycle matching can be combined with any distance.
An important downside to cycle matching is the need for resampling. In an exploratory context, the true data distribution is not known. Instead, Reani and Bobrowski [64] suggest to either bootstrap the existing finite dataset or sample from a kernel density estimate. Unfortunately, for high-dimensional data both approaches are problematic. Bootstrapping high-dimensional data is error-prone [18] and persistent homology of bootstraps is biased because repeated points effectively decrease the sample size [71]. Kernel density estimation requires a prohibitively large sample size in high dimensions [86, chapter 6.5].
Nevertheless, as a proof of concept, we explored cycle matching with the relatively fast ripser-based implementation of García-Redondo et al. [30]. The authors report a runtime of about $90$ minutes for a dataset of sample size $n=1000$ (their Table 3) and in our experiments we experienced even longer run times. As the result, this implementation of cycle matching was about $450$ times slower than our approach (Table S5). For this reason, we only used three resamples.
García-Redondo et al. [30]’s implementation accepts datasets as point clouds and implicitly assumes the Euclidean distance. Therefore, in this experiment we only used Euclidean distances and diffusion distances, which can be realized as Euclidean distances (Eq. 4). We computed prevalences of matched cycles on the noisy circle with $n=1000$ points in $\mathbb{R}^{50}$ for Gaussian noise of standard deviations $\sigma\in[0.1,0.2,0.25,0.3,0.35]$.
Figure S4: Results for cycle matching with the Euclidean distance on a noisy circle in $\mathbb{R}^{50}$ with noise level $\sigma$. Top row: Prevalences of the 10 most prevalent cycles. Means and standard deviation over three seeds. Color indicates whether the cycle was matched for all three random seeds or not. Second row: Persistence diagrams with the most prevalent features highlighted. Third row: Representative of the most prevalent feature overlaid on a 2D PCA of the data.
Figure S5: Results for cycle matching with the diffusion distance ($k=15,t=8$) on a noisy circle in $\mathbb{R}^{50}$ with noise level $\sigma$. Top row: Prevalences of the 10 most prevalent cycles. Means and standard deviation over three seeds. Color indicates whether the cycle was matched for all three random seeds or not. Second row: Persistence diagrams with the most prevalent features highlighted. Third row: Representative of the most prevalent feature overlaid on a 2D PCA of the data.
Cycle matching with the Euclidean distance did not reliably identify the ground-truth feature for $sigma>0.1$ (Figure S4). In all experiments with the Euclidean distance and $\sigma>0.1$, the feature with the highest prevalence did not correspond to the ground-truth loop.
For $\sigma=0.2,0.25$, we also checked the prevalences of the ground truth cycle. For $\sigma=0.2$ the ground truth cycle was not matched for any of the three resamples, leading to a prevalence of 00. For $\sigma=0.25$ the ground-truth cycle had the average prevalence of $0.1\pm 0.1$, much lower than the cycle with the highest prevalence.
One might hope that the cycles that do get matched for all resamples are true topological features, even if their prevalence values are low. Indeed, for $\sigma=0.1$ only the correct cycle was matched for all three resamples. However this heuristic did not work either because for $\sigma\in\{0.2,0.25,0.3\}$ no cycle was matched for all three resamples, while for $\sigma=0.35$ there was a cycle matched for all three resamples (not shown in Figure S4 because it was not among the 10 most prevalent ones), but it did not encode the ground-truth feature (and its prevalence $0.1\pm 0.1$ was very low).
The fact that most cycles were not matched for all resamples led to high standard deviations of the prevalences. The reason for the high mean prevalences for $\sigma>0.1$ was that some noise cycles got matched with very high prevalence (sometimes $>0.9$) for one or two of the resamples. For the same reason, the prevalence of the ground truth cycle for $\sigma=0.1$ was not much higher than the second highest prevalence. Increasing the number of resamples may help, but would make the procedure even more computationally expensive.
Overall, we conclude that cycle matching is not only very slow, but also does not alleviate the curse of dimensionality for persistent homology. In contrast, our spectral methods produced a near-perfect loop detection score until $\sigma=0.3$ (Figure 6).
That said, cycle matching can serve as an alternative performance score. We applied cycle matching to the diffusion distance (Figure S5) and found that
  1. 1.
    the maximal prevalences were larger than for the Euclidean distance for $\sigma<0.3$,
  2. 2.
    the cycle with maximal prevalence always represented the ground truth loop,
  3. 3.
    the prevalence of the correct cycle was a clear outlier among all prevalences for $\sigma\leq 0.3$,
  4. 4.
    the cycle representing the ground truth feature was always matched for all three resamples and was the only one matched for all three resamples for $\sigma\leq 0.3$.
This confirms the superior performance of spectral methods for detecting topology in high-dimensional settings.
Appendix F Effective resistance as a spectral method
In this section, we derive an explicit spectral embedding realizing the square root of von Luxburg et al. [81]’s corrected effective resistance. We also show that it is not necessarily a proper metric.
@@ Appendix F Effective resistance as a spectral method
Let us recall the notation from Sections 5 and 6, extended to weighted graphs. Let $G$ be a weighted connected graph with $n$ nodes. We denote by $A=(a_{ij})_{i,j=1,\dots,n}$ the weighted adjacency matrix whose entries $a_{ij}=a_{ji}$ equal the edge weight of edge $ij$ if edge $ij$ is part of the graph and zero otherwise. We further denote by $D=\text{diag}(d_{i})$ the degree matrix, where $d_{i}=\sum_{j}a_{ij}$ are the node degrees. We define $\text{vol}(G)=\sum_{i}d_{i}$. Let further $A^{\text{sym}}=D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$ and $L^{\text{sym}}=I-A^{\text{sym}}$ be the symmetrically normalized adjacency matrix and the symmetrically normalized graph Laplacian. Denote the eigenvectors of $L^{\text{sym}}$ by $u_{1},\dots,u_{n}$ and their eigenvalues by $\mu_{1},\dots,\mu_{n}$ in increasing order. The eigenvectors and eigenvalues of $A^{\text{sym}}$ are $u_{1},\dots,u_{n}$ and $1-\mu_{1},\dots,1-\mu_{n}$.
Definition F.1.
@@
$\frac{1}{\text{vol}(G)}H_{ij}=\frac{1}{d_{j}}+\langle b_{j},A^{\text{sym}}(b_{j}-b_{i})\rangle+\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle A^{\text{sym}}b_{j},u_{r}u_{r}^{\top}A^{\text{sym}}(b_{j}-b_{i})\rangle,$ (21)
where $b_{i}=\frac{1}{\sqrt{d_{i}}}\hat{e}_{i}=D^{-\frac{1}{2}}\hat{e}_{i}$ with $\hat{e}_{i}$ being the $i$-th standard basis vector.
@@
$\quad+\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle A^{\text{sym}}b_{j},u_{r}u_{r}^{\top}A^{\text{sym}}(b_{j}-b_{i})\rangle$ (24)
$\quad+\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle A^{\text{sym}}b_{i},u_{r}u_{r}^{\top}A^{\text{sym}}(b_{i}-b_{j})\rangle$ (25)
@@
$\quad+\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle A^{\text{sym}}(b_{j}-b_{i}),u_{r}u_{r}^{\top}A^{\text{sym}}(b_{j}-b_{i})\rangle$ (27)
For ease of exposition, we treat the two terms separately. By unpacking the definitions and using the symmetry of $D$, we get
@@
$\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle A^{\text{sym}}(b_{j}-b_{i}),u_{r}u_{r}^{\top}A^{\text{sym}}(b_{j}-b_{i})\rangle$ (31)
$=$ $\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle b_{j}-b_{i},(A^{\text{sym}}u_{r})(A^{\text{sym}}u_{r})^{\top}(b_{j}-b_{i}))\rangle$ (32)
$=$ $\sum_{r=2}^{n}\frac{1}{\mu_{r}}\langle b_{j}-b_{i},(1-\mu_{r})u_{r}\big{(}(1-\mu_{r})u_{r}\big{)}^{\top}(b_{j}-b_{i})\rangle$ (33)
@@
The corrected version of effective resistance and diffusion distances are, in general, not proper metrics, unlike the naive effective resistance [32, Corollary 2.4.]. We show this by giving concrete examples of graphs where the metric axioms for these distances do not hold.
Proposition F.3.
Neither corrected effective resistance nor diffusion distances between distinct points are necessarily positive. Moreover, corrected effective resistance does not always satisfy the triangle inequality.
@@
Consider the unweighted chain graph with three nodes (Figure S6, left). The uncorrected effective resistance between the first and the last node is $2$. So the corrected effective resistance between these distinct nodes is $2-1-1=0$. On the same graph, the random walker is necessarily at node $1$ after one step, independent whether it started at node 00 or $2$, so the diffusion distance with $t=1$ between these two distinct nodes is zero.
Consider now an unweighted graph with $5$ nodes, the first three of which form a triangle and the other two a chain connected to the triangle (Figure S6, right). Then
@@
We see that the triangle $0,2,4$ violates the triangle inequality
@@
The square root of corrected effective resistance does satisfy the triangle inequality and can also be used as input to persistent homology. This amounts to taking the square root of all birth and death times found with the corrected effective resistance as computing persistent homology commutes with strictly monotonic maps, and hence does not strongly affect the loop detection performance of corrected effective resistance (Figure S20). We therefore believe that it is not a problem for our persistent homology application that corrected effective resistance fails to satisfy the triangle inequality.
@@
Figure S6: Counterexample graphs for Proposition F.3
Appendix G Effective resistance integrates diffusion distances
First, we describe the connection between diffusion distances and naive effective resistance communicated to us by Mémoli [54]. To the best of our abilities, we could not find another reference. We use the same notation as in Section 6 and Appendix F.
Below, we speak of diffusion distances at half-integer time points, although the random walk analogy does not extend to half-steps. What we mean is inserting a half-integer value of $t$ in Eq. (4).
Proposition G.1.